OpenStudy (lgbasallote):

LGBARIDDLE!!! (you missed it didn't you ;D ) If: 72835 = 2 38562 = 3 10763 = 2 27321 = 0 97431 = 1 90078 = 5 23407 = 1 77412 = 0 33517 = 0 what is 95610?

OpenStudy (lgbasallote):

any ideas? :)

OpenStudy (anonymous):

Thinking..

OpenStudy (anonymous):

6

OpenStudy (lgbasallote):

hehe sure :) im a bit lagged up though :/

OpenStudy (anonymous):

5.5

OpenStudy (lgbasallote):

6 is wrong sorry

OpenStudy (lgbasallote):

btw show some explanation..dont guess lol

OpenStudy (anonymous):

but still by seeing the pattern..... 6 or 5.5 i think

OpenStudy (lgbasallote):

nope

OpenStudy (anonymous):

I think I got...

OpenStudy (lgbasallote):

shoot

OpenStudy (anonymous):

He he he.. That is working for only upper two.. Sorry...

OpenStudy (lgbasallote):

lol

OpenStudy (anonymous):

hmm

jhonyy9 (jhonyy9):

opinion 77412=7+7+4+1+2=21=0 so than 95610=9+5+6+1+0=21=0

OpenStudy (anonymous):

@jhonyy9 , 97431 = 1 90078 = 5 Their digit sum is also the same ie 24 but the result is not same..

OpenStudy (lgbasallote):

nope. it's not correct :(

jhonyy9 (jhonyy9):

yes ,true for this i have wrote this is just an opinion ,idea ...

OpenStudy (lgbasallote):

no no...no vectors involved...or matrices...this is merely an lgbariddle

OpenStudy (lgbasallote):

just look closely at the numbersd/digits

jhonyy9 (jhonyy9):

...2=3 ...3=2 ....1=0 95610=1

OpenStudy (anonymous):

No, that will also not work..

OpenStudy (lgbasallote):

hmm still no

OpenStudy (anonymous):

How do you do this?

OpenStudy (anonymous):

Is it just patterns?

OpenStudy (lgbasallote):

yup there's a pattern

jhonyy9 (jhonyy9):

72835=2 --- 38562=3 --- 3-8=-5+5=0+6=6/2=3 10763=2 --- 1*0=0*7=0+6=6/3=2 27321=0 --- 2+7=9/3=3-2=1-1=0 97431=1 --- 9-7=2+4=6/3=2-1=1 90078=5 --- 23407=1 --- 77412=0 --- 7/7=1-4=-3+1=-2+2=0 33517=0 --- 3+3=6-5=1-1=0*7=0 95610=0 --- 9-5=4-6=-2+1=-1*0=0

jhonyy9 (jhonyy9):

@lgbasallote opinion please !!!

OpenStudy (anonymous):

3.

OpenStudy (anonymous):

Should I disclose the explanation?

OpenStudy (anonymous):

@jhonyy9, your "reasoning" is flawed. Many of your "equations" are not true. e.g. 3-8=-5+5.

OpenStudy (anonymous):

i think the correct answer is 7 because certain patterns appear with the min/max element of each group after we add up the digits of each number,tell me if this is correct so i can further proceed with my explanation

OpenStudy (lgbasallote):

actually @Limitless is right..shall he disclose the explanation? haha

OpenStudy (anonymous):

@lgbasallote, we need spoiler tags or something. I'll PM you and you can medal me for it, I guess.